OpenStudy (anonymous):

Earn a medal What is the equation of the line passing through the points (–2, 2) and (3, –1)?

OpenStudy (anonymous):

first find the slope between these two points.

OpenStudy (anonymous):

slope is 3

OpenStudy (anonymous):

no! slope = y2 - y1 / x2 - x1 y2 - y1 = -1 - 2 = -3 x2 - x1 = 3 +2 = 5 so what is slope?

OpenStudy (anonymous):

how did gou find 3?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

NO! i said you try again! -3 / 5 is 4?!!!! :| :|

OpenStudy (anonymous):

do u know slope formula?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

i just told you.if you don't know slope formula then how do you find these answers? :D

OpenStudy (anonymous):

yes slope is: \[Slope = \frac{ y _{2} - y _{1} }{ x _{2} - x _{1} } \]

OpenStudy (anonymous):

now find slope.

OpenStudy (anonymous):

Find the slope (m), then just insert the numbers to this formula to find the equation: \[y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

@Dequavia123 , u still need help?

OpenStudy (anonymous):

besides this,u can find y-intercept and then put them in this equation: y = mx + d m -> slope d -> y-intercept